OpenStudy (anonymous):

Prove the following identity: (cos x + cos y)^2 + (sin x – sin y)^2 = 2 + 2cos(x + y)

OpenStudy (anonymous):

@dtan5457 Can you help me or find someone you can?

OpenStudy (anonymous):

@BreanneAndrews Can you help me or find someone you can?

OpenStudy (lυἶცἶ0210):

Try distributing the left hand side :)

OpenStudy (anonymous):

OKay

OpenStudy (anonymous):

Wrote the question down wrong...the two parentheses were supposed be squared. It should be fixed now @lυἶცἶ0210

OpenStudy (lυἶცἶ0210):

Yea, but distributing is still the first step :P Reply what you get when you distbute and I will help you from there ~

OpenStudy (anonymous):

okay, one sec

OpenStudy (anonymous):

I'm not sure if I did this write but this is what I got: \[\cos^2x + \cos^2y + \sin^2x - \sin^2y\]

OpenStudy (anonymous):

Are you still there?

OpenStudy (lυἶცἶ0210):

Well you got 2/3's of it, when you distribute: \(\ (cosx+cosy)^2+(sinx-siny)^2\) = \([(cosx*cosx)+(cosx*cosy)+(cosy*cosy)]\) + \([(sinx*sinx)+(sinx*-siny)+(-siny*-siny) ]\) (cut off so had to break up) But simplifing: \(\large = [(cos^2x+cosxcoxy +cos^2y)]+[(sin^2 x-sinxsiny +sin^2 y )\) Make sense so far?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

What's next?

OpenStudy (lυἶცἶ0210):

Use trig identities and group things: \(\large (sin^2x+cos^2x)+(sin^2y+cos^2y)+(cosxcosy-sinxsiny) \) Do you know your trig identities?

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

Don't most of the trig identities simplify things down to cos x or sin x? So wouldn't that them pointless in this situation since they're already simplified to cos x and sin x?

OpenStudy (lυἶცἶ0210):

By trig identities I mean ones like the Pythagorean Identities, odd-even identities, double-angle, sum and difference, etc.

OpenStudy (anonymous):

Oh yeah, I have a sheet next to me that has those. Do I use them now?

OpenStudy (lυἶცἶ0210):

Yup, use those to try and simplify this: \(\large (sin^2x+cos^2x)+(sin^2y+cos^2y)+(cosxcosy-sinxsiny)\)

OpenStudy (anonymous):

kk

OpenStudy (lυἶცἶ0210):

What did you end up with?

OpenStudy (anonymous):

\[((1 - \cos2x) / 2 + ((1 + \cos2x) / 2) + ((1 - \cos2x) / 2 + ((1 + \cos2x) / 2) + 1/2 (\cos (x - y) + \cos (x +y) - 1/2 (\cos (x - y) - \cos (x + y)\]

OpenStudy (lυἶცἶ0210):

Hmmm, try using the Pythagorean Identities :)

OpenStudy (anonymous):

One sec...

OpenStudy (anonymous):

I got 1 + 1 + (cos x cos y - sin x sin y) but I don't know what to do with the last part

OpenStudy (lυἶცἶ0210):

Right :) And the last identity isn't as common as the others, but you'll have to use the Sum and Difference Identities: http://prntscr.com/76van2

OpenStudy (lυἶცἶ0210):

So in the end it'll be 2+cos(x+y) which equals the right hand side :P

OpenStudy (anonymous):

Well actually the right side says 2 + 2 cos (x + y) but we only have one of the twos

OpenStudy (lυἶცἶ0210):

I feel silly now, I messed up in distributing in the beginning /.< It should of been: \(\large [(cos^2x+2cosxcoxy +cos^2y)]+[(sin^2 x-2sinxsiny +sin^2 y \) \(\large (sin^2x+cos^2x)+(sin^2y+cos^2y)+2(cosxcosy-sinxsiny) \) Now it'll simplify to 2+2cos(x+y) :P

OpenStudy (anonymous):

Okay, thanks! Do you mind if I ask you another question?

OpenStudy (lυἶცἶ0210):

Not at all, go ahead.

OpenStudy (anonymous):

Okay, here it is: Derive the trigonometric addition formula for sine: sin(a + b) = sin a cos b + cos a sin b

OpenStudy (lυἶცἶ0210):

Maybe you should make another post and see if someone else can help, I'm a bit stumped, sorry :/

OpenStudy (anonymous):

Okay, thanks for your help!